Find the z-score for the given raw score, mean, and standard deviation. Assume a normal probability distribution.
Raw score = 15, μ = 60, and o=12

Answer by drk(1908)

Z-score is x - MU/ SD or
Z -score = (15-60)/12 = -45/12 = - 3.75.
Using this you find percentages.